OpenStudy (anonymous):

Which of the following describes the correct process for solving the equation 3x + 6 = 21 and arrives at the correct solution? Divide both sides by 6 and then add 3. The solution is x = 5. Add 6 to both sides of the equation and then divide by 3. The solution is x = 9. Subtract 6 from both sides of the equation and then divide by 3. The solution is x = 8. Subtract 6 from both sides of the equation and then divide by 3. The solution is x = 5.

OpenStudy (anonymous):

@Baseballa101 @hyntc i just need help on 3 more problem please help

OpenStudy (anonymous):

D

OpenStudy (anonymous):

Emma took 35 pictures of flowers on a nature walk and an additional 3 pictures of each bird she saw. She took a total of 83 pictures. Which equation can be used to find the number of birds, x, that Emma saw on her walk? 35x + 3x = 83 35 + 3 = 83x 35x + 3 = 83 35 + 3x = 83

OpenStudy (anonymous):

help on this one too

OpenStudy (anonymous):

More??

OpenStudy (anonymous):

and then after this i just need help on 1 more prob. to see if i got it right @Baseballa101

OpenStudy (anonymous):

Ok